# Geometry of exactness of moment-SOS relaxations for polynomial optimization

### v3

@@ 5.1 Finite set
${\mathscr{S}}(2,2)^{1}_{1}=\mathbb{R}\oplus\{(f_{1},f_{2})\in\mathbb{R}^{2}:\left(\begin{array}[]{ccc}-4f_{1}-4f_{2}&\star&\star\\ f_{1}+s_{2}&-2s_{2}&\star\\ f_{2}-2s_{1}-s_{2}&s_{1}&4s_{1}+2s_{2}\end{array}\right)\succeq 0,(s_{1},s_{2})\in\mathbb{R}^{2}\}$
@@
On the left of Figure 1 we represent the first ($r=1$) moment relaxation ${\mathscr{R}}({\bf g})^{1}_{1}$ (dark gray), the convex hull $\mathrm{conv}\>{\mathscr{X}}$ (light gray), and the four points of ${\mathscr{X}}$ (black). The tiny dark gray region which remains visible are points in ${\mathscr{R}}({\bf g})^{1}_{1}\setminus\mathrm{conv}\>{\mathscr{X}}$. Also represented (in color) are the four normal cones at the four points. According to Lemma 2, up to the sign, they are the four spectrahedral cones ${\mathscr{S}}(\hat{{\bf x}})$, $\hat{{\bf x}}\in{\mathscr{X}}$ of Theorem 1. On the right of Figure 1 we represent the exactness cone ${\mathscr{F}}({\bf g})^{1}_{1}$ which is the union of the four spectrahedra, according to Theorem 1. We observe tiny conic regions (in white) corresponding to $\mathbb{R}[{\bf x}]_{1}\setminus{\mathscr{F}}({\bf g})^{1}_{1}$, namely first degree polynomials $f$ for which the moment-SOS relaxation of first order is not exact. If we solve the relaxation, we hit the slightly inflated tiny regions (dark gray on the left figure) of the moment relaxation ${\mathscr{R}}({\bf g})^{1}_{1}$, yielding a strict lower bound on the value $v(f)$. If instead we minimize the polynomials in ${\mathscr{F}}({\bf g})^{1}_{1}$, we hit one of the four points of ${\mathscr{X}}$, i.e. the relaxation is exact.
@@ 5.2 Non-convex set
On the left of Figure 3 we represent the first ($r=1$) moment relaxation ${\mathscr{R}}({\bf g})^{1}_{1}$ (dark gray) of ${\mathscr{X}}$ (light gray), as well as the normal cones to the points of the boundary of ${\mathscr{R}}({\bf g})^{1}_{1}$ where the first relaxation is exact. The green region is the normal cone to the left corner point of ${\mathscr{X}}$, the yellow region is the normal cone to the right corner point of ${\mathscr{X}}$, the blue line is the one-dimensional normal cone to the top corner point of ${\mathscr{X}}$, and the red region is the union of all the one-dimensional normal cones to the convex circular bottom part of ${\mathscr{X}}$. According to Lemma 2, up to the sign, the green, yellow, and blue cones are spectrahedral cones, whereas the red region is the union of spectrahedral cones along the circular arc. On the right of Figure 3 we represent the exactness cone ${\mathscr{F}}({\bf g})^{1}_{1}$ which is the union of these spectrahedral cones, according to Theorem 1. The blue line corresponds to the objective function $f({\bf x})=-x_{1}-x_{2}$ for which the first moment relaxation is exact. It is surrounded by a white region corresponding to objective functions for which the first moment relaxation is not exact. The other colored regions belong to the exactness cone.
@@ Acknowledgement
The statement of Theorem 1 and its proof were given by an anonymous reviewer.

### v4

@@ 5 Examples
Data availability statement: The Matlab codes to generate the figures in the following examples are available upon request.
@@ 5.1 Finite set
${\mathscr{S}}(2,2)^{1}_{1}=\mathbb{R}\oplus\{(f_{1},f_{2})\in\mathbb{R}^{2}:\left(\begin{array}[]{ccc}-4f_{1}-4f_{2}&{\color[rgb]{0,0,0}f_{1}+s_{2}}&{\color[rgb]{0,0,0}f_{2}-2s_{1}-s_{2}}\\ f_{1}+s_{2}&-2s_{2}&{\color[rgb]{0,0,0}s_{1}}\\ f_{2}-2s_{1}-s_{2}&s_{1}&4s_{1}+2s_{2}\end{array}\right)\succeq 0,(s_{1},s_{2})\in\mathbb{R}^{2}\}$
@@
On the left of Figure 1 we represent the first ($r=1$) moment relaxation ${\mathscr{R}}({\bf g})^{1}_{1}$ (dark gray), the convex hull $\mathrm{conv}\>{\mathscr{X}}$ (light gray), and the four points of ${\mathscr{X}}$ (black). The tiny dark gray region which remains visible are points in ${\mathscr{R}}({\bf g})^{1}_{1}\setminus\mathrm{conv}\>{\mathscr{X}}$. Also represented (in color) are projections on the coordinates $(f_{1},f_{2})$ of the four normal cones at the four points. According to Theorem 2, up to the sign, they are the four spectrahedral cones ${\mathscr{S}}(\hat{{\bf x}})$, $\hat{{\bf x}}\in{\mathscr{X}}$ of Theorem 1. On the right of Figure 1 we represent the exactness cone ${\mathscr{F}}({\bf g})^{1}_{1}$ which is the union of the four spectrahedra, according to Theorem 1. We observe tiny conic regions (in white) corresponding to $\mathbb{R}[{\bf x}]_{1}\setminus{\mathscr{F}}({\bf g})^{1}_{1}$, namely first degree polynomials $f$ for which the moment-SOS relaxation of first order is not exact. If we solve the relaxation, we hit the slightly inflated tiny regions (dark gray on the left figure) of the moment relaxation ${\mathscr{R}}({\bf g})^{1}_{1}$, yielding a strict lower bound on the value $v(f)$. If instead we minimize the polynomials in ${\mathscr{F}}({\bf g})^{1}_{1}$, we hit one of the four points of ${\mathscr{X}}$, i.e. the relaxation is exact.
@@ 5.2 Non-convex set
On the left of Figure 3 we represent the first ($r=1$) moment relaxation ${\mathscr{R}}({\bf g})^{1}_{1}$ (dark gray) of ${\mathscr{X}}$ (light gray), as well as the normal cones to the points of the boundary of ${\mathscr{R}}({\bf g})^{1}_{1}$ where the first relaxation is exact. The green region is the normal cone to the left corner point of ${\mathscr{X}}$, the yellow region is the normal cone to the right corner point of ${\mathscr{X}}$, the blue line is the one-dimensional normal cone to the top corner point of ${\mathscr{X}}$, and the red region is the union of all the one-dimensional normal cones to the convex circular bottom part of ${\mathscr{X}}$. According to Theorem 2, up to the sign, the green, yellow, and blue cones are spectrahedral cones, whereas the red region is the union of spectrahedral cones along the circular arc. On the right of Figure 3 we represent the exactness cone ${\mathscr{F}}({\bf g})^{1}_{1}$ which is the union of these spectrahedral cones, according to Theorem 1. The blue line corresponds to the objective function $f({\bf x})=-x_{1}-x_{2}$ for which the first moment relaxation is exact. It is surrounded by a white region corresponding to objective functions for which the first moment relaxation is not exact. The other colored regions belong to the exactness cone.
@@ Acknowledgement
The proofs of Theorems 1 and 2, as well as a simplified statement of Definition 1, were suggested by anonymous reviewers. I am also grateful to David de Laat for useful comments.
Conflict of interests
I do not have any conflict of interest to declare regarding this submission.